(2500/0.07)(1-(1+0.05)^10. I do I set up an exponent on a scientific calculator
Answer by RAY100(1608)

(2500/.07)(1-(1+.05)^10)
(35,714.29)(1-(1.05)^10)
35,714.29 (1-1.629)
35,714.29(-.629)
-22,464.29,,,,,,,,,(-22,460.52 without rounding)
Note, there are 2 kind of exponent keys on scientific calculators.
On Texas instrument calc,,in the right column,above the divide key, pls find a key labeled,
(y^x),,,it actually shows the x just raised without carrot sign.
to use try 3^2, just strike 3 key then (y^x) key then 2 key then equal key,,should be 9
use this same key, but in 2nd level to find square root.
try sq rt 9, strike 9 key, then "2nd" key, then (y^x) key, then 2 key , then equal key,,,,should be 3.
Other calculators use same logic but use a "^" key instead of the "y^x" key.
Sometimes these are located almost in the center of the calculator.
you might try this on some comon squares, remember 2^2=4, 3^2+9, 4^2=16,etc
on very questionable calculations, remember that raising to power is just how many times a number is multiplied by itself, for example 2^3=2*2*2=8, or 2^4=2*2*2*2=16
Good Luck and keep up good work